OpenStudy (anonymous):

Solve using the substitution method . x + y = 1 5x-4y=-7

OpenStudy (anonymous):

x=1-y 5-5y-4y=-7 -9y=-12 y=12/9 y=4/3

OpenStudy (anonymous):

medal please